OpenStudy (legomyego180):

Determine whether the series converges or diverges:

OpenStudy (legomyego180):

\[\sum_{}^{}\frac{ k+1 }{ \sqrt{2k^3+k} }\]

OpenStudy (legomyego180):

@ganeshie8

OpenStudy (legomyego180):

I think I need to use limit comparison test?

OpenStudy (phi):

yes, but which bn ?

OpenStudy (legomyego180):

This is a picture of how one of my classmates did the question, but I dont understand some of his steps.

OpenStudy (legomyego180):

\[\frac{ 1 }{ k^3 }\] ? I say that because I think you usually want to pick the highest power right?

OpenStudy (evoker):

Your classmate divided each term by k.

OpenStudy (legomyego180):

why?

OpenStudy (legomyego180):

and \[\frac{ 2k^3 }{ k } \neq 2k\] So that part in the denominator is confusing me too

OpenStudy (phi):

if you divide all terms by k you get \[ \frac{1+\frac{1}{k}}{\sqrt{2k + \frac{1}{k}} }\] maybe consider 1/ sqrt(2k) ? we know 1/k diverges

OpenStudy (legomyego180):

Sorry if im stuck on the basics here. is the denominator \[\sqrt{2k+\frac{ 1 }{ k }}\] because \[\sqrt{2k^3+k}= \sqrt{2k(k^2+1)}\]? Im not sure how he divided the stuff in the radical by k :/

OpenStudy (phi):

use k= sqrt(k^2)

OpenStudy (legomyego180):

oh oh course

OpenStudy (phi):

so \[ \frac{\sqrt{2k^3+k}}{k} = \sqrt{ \frac{2k^3+k}{k^2} }\]

OpenStudy (phi):

I think it's clear it diverges. Just need the details

OpenStudy (legomyego180):

Yea. Can I use basic divergence test at this point?

OpenStudy (legomyego180):

Or do I need to use P Series

ganeshie8 (ganeshie8):

See if below manipulation looks any simpler : \[\frac{ k+1 }{ \sqrt{2k^3+k} } \ge\frac{ k+1 }{ \sqrt{2k^3+2k^3} } \ge \dfrac{k}{2k\sqrt{k}} = \dfrac{1}{2\sqrt{k}} \]

OpenStudy (legomyego180):

That makes sense @ganeshie8

ganeshie8 (ganeshie8):

See if below manipulation looks any simpler : \[\frac{ k+1 }{ \sqrt{2k^3+k} } \ge\frac{ k+1 }{ \sqrt{2k^3+2k^3} } = \dfrac{k+1}{\sqrt{4k^3}} \ge \dfrac{k}{2k\sqrt{k}} = \dfrac{1}{2\sqrt{k}} \]

OpenStudy (legomyego180):

So really the trick here is just rewriting using algebra and then using p series

OpenStudy (phi):

It looks like the trick is to do a bunch of these and learn the techniques. Otherwise, you will be like me and wander about in search of an idea.

OpenStudy (legomyego180):

We are in the same Boat @phi lol

OpenStudy (legomyego180):

@ganeshie8 did you replace k with 2k^3 just so it would be easy to pull out of the radical?

ganeshie8 (ganeshie8):

Yes, but why are we allowed to replace like that ?

OpenStudy (welshfella):

Yes - In college I learned about 20 different tests for convergence but I cant access them in my brain anymore!!

OpenStudy (legomyego180):

I dont know. I thought you could if the values converged, but this diverges.

ganeshie8 (ganeshie8):

Haha there are more than a dozen convergence tests. I remember only the comparison test ..

ganeshie8 (ganeshie8):

And the geometric series, ofcourse !

OpenStudy (welshfella):

yea

OpenStudy (legomyego180):

Yea, Im really hoping I dont have to remember them next semester in Calc 3

ganeshie8 (ganeshie8):

Let me break it into few steps for you

ganeshie8 (ganeshie8):

Would you agree that \(k\le 2k^3\) is true for all \(k\gt 1\) ?

OpenStudy (legomyego180):

yea for sure

ganeshie8 (ganeshie8):

Add \(2k^3\) both sides and get \(k + \color{red}{2k^3} \le 2k^3+ \color{red}{2k^3} \)

ganeshie8 (ganeshie8):

Still with me ?

OpenStudy (legomyego180):

yes

ganeshie8 (ganeshie8):

Take square root both sides and get \(\sqrt{k + \color{red}{2k^3} }\le \sqrt{2k^3+ \color{red}{2k^3}} \)

ganeshie8 (ganeshie8):

Next, take flip both sides. What happens to the inequality when you flip ?

OpenStudy (legomyego180):

It also flips

ganeshie8 (ganeshie8):

@legomyego180 Calc3 is mostly about line integrals and surface integrals. But it helps to remember comparison test and geometric series. Some proofs in Calc3 may use these.

ganeshie8 (ganeshie8):

Flipping both sides gives \(\dfrac{1}{\sqrt{k + \color{red}{2k^3} }}\ge \dfrac{1}{\sqrt{2k^3+ \color{red}{2k^3}} }\)

ganeshie8 (ganeshie8):

Multiply k+1 both sides

OpenStudy (legomyego180):

I see now. Because \[\frac{ k+1 }{ \sqrt{2k^3+3} }\] is greater than \[\frac{ 1 }{ 2k^3+2k^3 }\] (forgot the square root sign) And because \[\frac{ 1 }{ \sqrt{2k^3+2k^3} }\] diverges that means our original equation must also diverge because it is greater.